OpenStudy (evansanchez):

Determine if the solution set for the system of equations shown is the empty set, contains one point or is infinite. 2x - y = 7 2y = 4x - 14 {} 1 solution infinite

OpenStudy (solomonzelman):

\(\color{black}{2x - y = 7\quad \Longleftrightarrow \quad 2x - y \color{red}{-2x}= 7\color{red}{-2x}\quad \Longleftrightarrow \quad - y = 7-2x}\) \((\color{black}{\quad \Longleftrightarrow \quad \color{red}{(-2)(}- y \color{red}{)}= \color{red}{(-2)(}7-2x\color{red}{)}\quad \Longleftrightarrow \quad 2y=-14+4x} \).

OpenStudy (solomonzelman):

Can you tell me anything about the equations?\(\\[0.5em]\) (How they relate, any differences/similarities?)

OpenStudy (solomonzelman):

Your first equation: 2y = 4x - 14 Your second equation: 2y = -14 + 4x

OpenStudy (solomonzelman):

Do you notice anything about the equations?